OpenStudy (anonymous):

Ellen works for a high-speed rail company that wants to develop a new rail line. Ellen’s project is to find a train that is the second fastest in the world. The Shinkansen Bullet Train in Japan is reported to go as fast as 320 kilometers per hour. The TGV train in France can reach speeds of 89.44 meters per second. Explain to Ellen how to find what speed her new train must go to be the second fastest in this group. Then, find an appropriate speed for Ellen’s train in miles per hour and use the formula distance = speed • time to find how far the train can go in 2.5 hours. Use complete sentenc

OpenStudy (anonymous):

Which part is giving you trouble?

OpenStudy (anonymous):

all of it

OpenStudy (anonymous):

1 hour=360 seconds (60*60) 1 kilometer=1000 meters to convert kilometers/hour to meters/second you can just do 1(kilometer)/1(hour)=1000(meters)/360(seconds) seeing as how you want your final answer in hours well convert the 89.44m/s train into km/h 89.44*1000/360=?km/h then you just pick a speed somewhere in between the two numbers and convert it into miles per hour (mph) 1mph=1.60934km/h so x/1.60934=?mph for the 2.5 hours distance traveled you can just plug in the time (2.5 hours) and your chosen speed into the equation you listed above which will give you how many miles the train has travelled

OpenStudy (anonymous):

okay thats it